OpenStudy (anonymous):

Consider the line .y=-2/3 x+69 Find the equation of the line that is parallel to this line and passes through the point .(9,4) Find the equation of the line that is perpendicular to this line and passes through the point .(9,4)

OpenStudy (mattfeury):

can you tell me the slope of that line (in the equation)?

OpenStudy (anonymous):

-2/3

OpenStudy (anonymous):

hey i wrote it wrong

OpenStudy (anonymous):

its 6 not 69

OpenStudy (mattfeury):

oh ok. pretty similar still. so the equation for a line (including that one) is y = mx + b we know that m = -2/3. and we want a line that passes through (9,4), aka x=9, y=4. so we can solve for b.

OpenStudy (mattfeury):

\[y = mx+b\]\[4 = -\frac{2}{3}*9 + b \]

OpenStudy (mattfeury):

think ya can solve that?

OpenStudy (anonymous):

it is asking for a perpendicular and parallel so two different answers

OpenStudy (mattfeury):

right. this is just for the first part. it's a two part question.

OpenStudy (mattfeury):

this is the equation of the first line, but we need to solve for b to get it. we'll get to the second one in a sec :p

OpenStudy (anonymous):

aww ok so were finging the slope right now?

OpenStudy (anonymous):

b=y+6

OpenStudy (mattfeury):

oh no, sorry. i guessed i skipped over that, heh. when lines are parallel, their slopes are the same. so we have a first line with a certain slope, and we want a second line that will also have that same slope. they will be parallel (never intersecting). when lines are perpendicular, their slopes are "reciprocals," meaning flipped sorta. the reciprocal of 2 is 1/2 (because 2 = 2/1). the reciprocal of 3/4 is 4/3.

OpenStudy (mattfeury):

\[ -\frac{2}{3}*9 = -\frac{18}{3} = -6\]

OpenStudy (anonymous):

if they're perpendicular they have to be opposite reciprocals, so 2 would be -1/2

OpenStudy (mattfeury):

oops. good catch stacey! must be getting sleepy.

OpenStudy (anonymous):

man this is hard stuff i dont know how you guys can do this

OpenStudy (mattfeury):

it's not too bad once it clicks. you just need to stick with it and try not get frustrated. it becomes difficult to learn at that point. so we need a line. we know an equation for a line is: y = m*x + b. we need to know 'm' and 'b' though, in order to make it an actual equation. we know m is slope. and since this line is parallel to our other line, it has the same slope. now we just need to solve for b and we'll have an equation :)

OpenStudy (mattfeury):

lost? lol

OpenStudy (anonymous):

yes i am =/

OpenStudy (mattfeury):

any place in particular?

OpenStudy (mattfeury):

do you follow any of what i said just above^^

OpenStudy (anonymous):

-2/3*9 is that what i have to do to get the b?

OpenStudy (mattfeury):

sorta. don't jump too far ahead. do you understand where I got\[4 = -\frac{2}{3}*9 + b\]from?

OpenStudy (anonymous):

aww..ok so then if i follow through witht hat then b =10

OpenStudy (mattfeury):

yup :)

OpenStudy (mattfeury):

so substitute 'm' and 'b' into our equation for a line

OpenStudy (anonymous):

would m be 6

OpenStudy (mattfeury):

m is our slope, which we know is the same as the other line, -2/3

OpenStudy (anonymous):

so 4=-2/3*9+10

OpenStudy (mattfeury):

yea but you can leave y and x in there. the equation is y = -2/3*x + 10. this is an equation of a parallel line that passes through point (9,4). you can check and plug x=9, y=4 in that equation and it should be equal.

OpenStudy (anonymous):

aww well made that a little hard.. so then for a perp. line would it be the recipricle of -2/3

OpenStudy (anonymous):

i made that a little hard***

OpenStudy (mattfeury):

heh no problemo. as long as it makes sense!

OpenStudy (mattfeury):

and yeah, we would do the same thing, but this time m = 3/2, the negative reciprocal of -2/3.

OpenStudy (mattfeury):

that would give us the equation of a perpendicular line

OpenStudy (anonymous):

so y=3/2x+10

OpenStudy (mattfeury):

it's not necessarily 10 :p. you'll have to use 'b' and solve for b.

OpenStudy (anonymous):

so ill have to do the same equation but change the slope

OpenStudy (mattfeury):

no you need to find the y-intercept, b. you know the slope is 3/2.

OpenStudy (anonymous):

b=-19/2

OpenStudy (mattfeury):

i think you're right. :) let me double check. can you put the equation together now?

OpenStudy (anonymous):

yeah that helped out a lot ty

OpenStudy (mattfeury):

np :)

OpenStudy (mattfeury):

you're right, ps. nice job!

OpenStudy (anonymous):

sweet got a few more problems like this i think ill be able to do now!

OpenStudy (anonymous):

well i have another question if the slope is 7 the recipricle would be what?

OpenStudy (anonymous):

like -1/7